OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#13572 — Optical loop Paris
Incident Report for Web Cloud
Resolved
A fibre outage impacted the loop in Paris.
We've opened a ticket with our fibre operator.

Update(s):

Date: 2015-06-12 10:26:59 UTC
During the night, SFR welded the fibre to patch the outage.
The loop is now secure.

Date: 2015-06-09 11:20:21 UTC
The fibre outage caused the network to behave abnormally. It's been fixed.
We will continue to monitor the network until the optical loop is secure.
Posted Jun 09, 2015 - 11:17 UTC